OpenStudy (anonymous):

What is 7.8 × 105 m/h in kilometers per day?

OpenStudy (anonymous):

A. 3.24 × 1010 km/day B. 3.24 × 108 km/day C. 1.872 × 106 km/day D. 1.872 × 104 km/day

OpenStudy (triciaal):

welcome to openstudy. do you see that you will need to do 2 sets of conversions m to km and h to day?

OpenStudy (anonymous):

first you know that 1 m = 1/1000 km and 1h = 1/24 day then you can apply these equation